OpenStudy (anonymous):

33 A.prime B.composite :1,33 C.composite:1,3,11,33 D.composite:1,3,6,11,33

OpenStudy (dayakar):

33= 1*33 = 3*11

OpenStudy (dayakar):

factors of 33 = {1,3,33}

OpenStudy (anonymous):

A,B,C,or D

OpenStudy (dayakar):

any number which has more than two factors is called composite number

OpenStudy (anonymous):

so what the answer

OpenStudy (dayakar):

c